OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y-6=-3x-3 y-6=-3(x+1) y=-3x+3 3x+y=3 @astrophsics

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

@pooja195 would it be -1,6

pooja195 (pooja195):

can you write it as the whole equation and plug in the given values?

OpenStudy (anonymous):

Wait it's not like the other question. I think it's B @pooja195

pooja195 (pooja195):

Good :D you got it! :)

OpenStudy (anonymous):

Oh okay. May I ask you another one @pooja195

pooja195 (pooja195):

Post a new question.